OpenStudy (anonymous):

steps involved in solving 2nd order differential equation

OpenStudy (anonymous):

depends on method used. Most simple case is 2º order ODE with constant coeficients. 1º step: find caracteristic polinomial 2º solve it for 0 3º write the solution

OpenStudy (anonymous):

thanks,I mean using power series method

OpenStudy (anonymous):

1º step : supose the solution is analytic. It means can be written as power series 2º step: find the expression of each of it's derivatives as power series 3º put that into the ODE 4º make equal indices 5º solve

OpenStudy (anonymous):

Thanks, what is ODE?

OpenStudy (anonymous):

ordinary differential equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve using simultaneous ordinary differential equation using elimination method dx/dt=x+y, dy/dt=4x-2y

OpenStudy (anonymous):

\[\huge \color{green}{\text{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

thanks 99

OpenStudy (anonymous):

waterineyes,how're you?

OpenStudy (unklerhaukus):

Ordinary differential equations are distinct from Partial differential equations , which have ODE's have PDE's have \(\frac{\text d y}{\text d x}\)'s \(\frac{\partial y}{\partial x}\)'s